OpenStudy (anonymous):

Which of the following equations does not require two steps to solve? negative two a minus ten equals nine eleven minus x equals 3 t minus fifteen equals negative seven negative eighteen plus two n equals negative six if you help i will give a medal and become a fan

OpenStudy (anonymous):

some one please help me

OpenStudy (anonymous):

Do you want the answer or do you want to think it through?

OpenStudy (anonymous):

i need the answer

OpenStudy (anonymous):

I see two of them that are one-step answers, the rest are multiple-step answers, so technically ALL of them do not require two answers.

OpenStudy (anonymous):

errr steps.

OpenStudy (anonymous):

let's first turn the words into numbers and variables. We have: -2a - 10 = 9 11 - x = 3 t - 15 = -7 -18 + 2 = -6

OpenStudy (anonymous):

On each one, the first thing you need to do is isolate the variable

OpenStudy (anonymous):

sigh i mis-typed the last equation. it should read -18 +2n = -6

OpenStudy (anonymous):

on how many of them can you isolate a positive variable in one step?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

okay, let's work b. subtract 11 from both sides.

OpenStudy (anonymous):

what are you left with?

OpenStudy (anonymous):

14 i think right

OpenStudy (anonymous):

11 - x = 3 11 - x - 11 = 3 - 11 -x = ?

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

correct. now let's make that negative variable positive.

OpenStudy (anonymous):

to that, we multiply both sides by -1

OpenStudy (anonymous):

what do you mean by that

OpenStudy (anonymous):

11 - x = 3 11 - x - 11 = 3 - 11 -x = -8 -1 * x = -8 * -1 x = 8 this is 2 steps because you first subtracted 11 from both sides, then multiplied by -1

OpenStudy (anonymous):

re-reading, there is 2 more that require at least 2 steps, and 1 that is one single step. i apologize for mis-typing earlier

OpenStudy (anonymous):

c

OpenStudy (anonymous):

very good. now show me why it only requires one step.

OpenStudy (anonymous):

because when you do the reverse operation you do 15+7=22 so that is how you get the answer

OpenStudy (anonymous):

first thing you want to do in all 4 of these questions is to isolate the variable. in this case, you simply need to add 15 to both sides. in this case, the answer is t - 15 = -7 t - 15 + 15 = -7 + 15 t = ?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

technically it is two steps because first you are adding something to both sides, then in second step you are combining like terms, but since combining like numbers is "automatic" it is not considered a step by itself

OpenStudy (anonymous):

but your instructor will most likely choose c as the "not two steps" fulfillment of the question

OpenStudy (anonymous):

thank you so much @Firebrand

OpenStudy (anonymous):

:) my pleasure! I hope this helps you with future problems similar to these